A machine to manufacture fasteners has a setup cost of $1,300 and a unit cost of $0.009 for each fastener manufactured. A newer machine has a setup cost of $1,700 but a unit cost of only $0.0025 for each fastener manufactured. Find the break point. (Round your answer to the nearest whole unit.)
Answer by checkley79(3341)

1300+.009X=1700+.0025X
.009X-.0025=1700-1300
.0065X=400
X=400/.0065
X=61,538 FASTENERS.
PROOF:
1300+.009*61,538=1700+.0025*61,538
1300+553.85=1700+153.85
1853.85=1853.85